OpenStudy (anonymous):

Solve the equation by factoring. 4a^2 - 16a = -15

OpenStudy (anonymous):

i dont know sorry :(

OpenStudy (anonymous):

thats fine

OpenStudy (anonymous):

ok whats next

OpenStudy (perl):

\[4a^2-16a =-15 \\ 4a^2-16a + 15 = 0 \\ 4a^2-6a - 10a + 15 = 0 \\ 2a(2a - 3) -5(2a - 3)=0 \\ (2a-5)(2a-3)=0\]

OpenStudy (anonymous):

what next

OpenStudy (perl):

now you can use zero product property, set the factors equal to zero and solve